# Cyclic- and helical-symmetry-adapted phonon formalism within density functional perturbation theory

### v1

@@ I Introduction
Low-dimensional materials have attracted increasing attention over recent decades due to their distinctive mechanical, electronic, vibrational, and thermal properties Xia et al. (2003). These material systems are not limited to engineered nanostructures such as nanosheets, nanotubes, nanorods, nanowires, nanoribbons, nanodots, and nanoclusters, but also occur naturally in the form of viruses, RNA, and proteins. Non-translational symmetries are ubiquitous in such systems, with cyclic and helical symmetries being among the most prevalent James (2006); Allen (2007a). Even when absent in the undeformed configurations, mechanical deformations such as bending and twisting can induce cyclic and helical symmetries James (2006); Kit et al. (2011). Exploiting these symmetries leads to substantial computational savings and simplifications in the analysis of physical properties, motivating the development of cyclic- and helical-symmetry-adapted formulations in atomistic force-field methods James (2006); Dumitrica and James (2007); Dayal and James (2010); Aghaei and Dayal (2011); Aghaei et al. (2013), tight-binding approaches White et al. (1993); Allen (2007b, c); Popov (2004); Popov and Lambin (2006); Gunlycke et al. (2008); Zhang et al. (2009a, b); Kit et al. (2011); Zhang and Wei (2017), machine-learned force fields (MLFFs) Sharma et al. (2025a), and Kohn–Sham DFT Saito et al. (1992); Ono and Hirose (2005a, b); Banerjee and Suryanarayana (2016); Ghosh et al. (2019); Banerjee (2021); Sharma and Suryanarayana (2021). In particular, the symmetry-adapted Kohn–Sham framework has enabled diverse applications, including bending moduli of 2D materials Kumar and Suryanarayana (2022, 2020), elastic moduli of nanotubes Bhardwaj et al. (2021); Bhardwaj and Suryanarayana (2022a), the flexoelectric effect in 2D materials Codony et al. (2021); Kumar et al. (2021), and impact of mechanical deformations on the electronic Bhardwaj and Suryanarayana (2023, 2022b) and spintronic Bhardwaj and Suryanarayana (2024) properties of nanotubes. However, these studies are restricted to static behavior and response.
@@ II Cyclic and helical symmetry
Consider a quasi-one-dimensional nanostructure comprising of $\hat{N}_{a}$ fundamental atoms and exhibiting cyclic and/or helical symmetry, as illustrated in Fig. 1. Let $\tilde{\theta}$ denote the angle between an atom and its nearest cyclic image, and let $\varphi$ denote the circumferential-plane angle between an atom and its nearest helical image, which is separated by a distance $H$ along the longitudinal direction. The symmetry group of the nanostructure can be represented as a set of isometries obtained from the direct product (denoted by $\times$) of a cyclic symmetry group $\mathfrak{C}$ and a helical symmetry group $\mathfrak{H}$ McWeeny (2002):
@@
Indeed, the nanostructure is periodic along both the $\tilde{\theta}$ and $z$ directions in this helical coordinate system. Therefore, the helical coordinates of the image atom $\mathbf{\tilde{R}}_{I^{\prime}}$ are related to those of the corresponding fundamental atom $\mathbf{\tilde{R}}_{I}$ by a translation vector, i.e., $\mathbf{\tilde{R}}_{I^{\prime}}=\mathbf{\tilde{R}}_{I}+(0\,\,\zeta_{I^{\prime}}\tilde{\Theta}\,\,\mu_{I^{\prime}}H)^{T}$. Moreover, the 2D reciprocal space associated with the nanostructure in this coordinate system is discrete-continuous and periodic, with the first Brillouin zone (denoted by $\mathcal{U}$) defined by the set of $|\mathscr{G}|$ points $(\nu,\eta)$ such that, $\nu\in\{0,1,2,\cdots,\mathfrak{N}-1\}$ and $\eta\in[-\pi/H,\pi/H)$. These Brillouin zone points are commonly referred to as wavevectors. In what follows, the electron and phonon wavevectors are denoted by $\mathbf{k}=(0,\nu_{\mathbf{k}},\eta_{\mathbf{k}})$ and $\mathbf{q}=(0,\nu_{\mathbf{q}},\eta_{\mathbf{q}})$, respectively.
@@ III Symmetry-adapted DFT
The Hamiltonian operators depend on the electron density, which in turn depends on the orbitals, and therefore the governing equations must be solved iteratively to achieve self-consistency. The resulting electronic ground-state quantities can then be used to evaluate the energy and its first-order derivatives, which includes the atomic forces and stress tensor Sharma and Suryanarayana (2021).
@@ IV.1 Perturbative expansion
In the limit of small perturbations, the atomic displacement in the Cartesian coordinate system can be obtained using the transformation:
@@ IV.2 Sternheimer equation
where $\mathcal{I}$ denotes the identity operator, and the coefficients $\xi_{m,\mathbf{k}+\mathbf{q}}$ are chosen to remove the singularity and improve the conditioning of the equation, thereby ensuring robust and faster convergence to the solution. .
@@ V Implementation
We have implemented the cyclic- and helical-symmetry-adapted Kohn-Sham DFT phonon framework within the Cyclix-DFT Sharma and Suryanarayana (2021) feature of the real-space code M-SPARC Xu et al. (2020); Zhang et al. (2023), which is a Matlab-based variant of the large-scale electronic-structure code SPARC Xu et al. (2021); Zhang et al. (2024). In particular, all quantities in the Kohn-Sham formalism are represented on a uniform grid in the aforementioned helical coordinate system. The differential operators are approximated using high order centered finite-differences and spatial integrations are approximated using the trapezoidal rule. The Brillouin zone is discretized at the points $(\nu,\eta_{MP}+\frac{\varphi}{H}\nu)$, where $\eta_{MP}$ is generated using the Monkhorst-Pack Monkhorst and Pack (1976) scheme. The Purdew-Zunger Perdew and Zunger (1981) variant of the LDA Kohn and Sham (1965) exchange-correlation, optimized norm-conserving Vanderbilt pseudopotentials Hamann (2013), and Fermi-Dirac electronic smearing are employed.
@@ VI Results and discussion
where $a=\sqrt{3}\,a_{0}$ is the length of graphene’s lattice vector and $r$ is the nanotube radius. This scaling law is in very good qualitative agreement with previous studies Sharma et al. (2025b); Gunlycke et al. (2008). The prefactor of $45~\mathrm{cm}^{-1}$ obtained here is is very good agreement with the value of $46~\mathrm{cm}^{-1}$ reported recently using a cyclic- and helical-symmetry-adapted machine-learned force field Sharma et al. (2025b), but differs noticeably from the value of $54~\mathrm{cm}^{-1}$ obtained earlier using an interatomic potential Gunlycke et al. (2008).
@@
This scaling law is also in very good qualitative agreement with the literature Sharma et al. (2025b); Gunlycke et al. (2008). However, the constant of $468$ cm-1 obtained here is noticeably different from the values of $480$ cm-1 and $486$ cm-1 obtained previously using machine-learned force field Sharma et al. (2025b) and interatomic potential Gunlycke et al. (2008), respectively.
@@ VII Concluding remarks
In this work, we developed a first-principles framework for the calculation of phonons in nanostructures with cyclic and/or helical symmetry. In particular, we derived cyclic- and helical-symmetry-adapted representations for the dynamical matrix at arbitrary phonon wavevectors within a variationally formulated, symmetry-adapted DFPT framework. In addition, we derived the acoustic sum rules for cylindrical geometries, which include a rigid-body rotational mode in addition to the three translational modes. The resulting formalism was implemented within a high-order finite-difference discretization. Using carbon nanotubes as representative systems, we validated the accuracy of the framework through excellent agreement with periodic plane-wave results. We further applied the framework to compute the Young’s and shear moduli of carbon nanotubes, as well as the scaling laws governing the dependence of ring and radial breathing mode phonon frequencies on nanotube diameter. The elastic moduli were found to be in agreement with previous DFT and experimental results, while the phonon scaling laws showed qualitative agreement with previous atomistic simulations.
@@ ACKNOWLEDGMENTS
A.S. gratefully acknowledges the ANRF early career research grant (Grant No. ANRF/ECRG/2024/002362) from Department of Science and Technology India. P.S. gratefully acknowledges the support of the U.S. National Science Foundation under grant CAREER–155321, and the U.S. Department of Energy, Office of Science under grant DE-SC0023445. This research was also supported by the supercomputing infrastructure provided by Partnership for an Advanced Computing Environment (PACE) through its Phoenix cluster at Georgia Institute of Technology, Atlanta, Georgia.

### v2

@@ I Introduction
Low-dimensional materials have attracted increasing attention over recent decades due to their distinctive mechanical, electronic, vibrational, and thermal properties Xia et al. (2003). These material systems are not limited to engineered nanostructures such as nanosheets, nanotubes, nanorods, nanowires, nanoribbons, nanodots, and nanoclusters, but also occur naturally in the form of viruses, RNA, and proteins. Non-translational symmetries are ubiquitous in such systems, with cyclic and helical symmetries being among the most prevalent James (2006); Allen (2007a). Even when absent in the undeformed configurations, mechanical deformations such as bending and twisting can induce cyclic and helical symmetries James (2006); Kit et al. (2011). Exploiting these symmetries leads to substantial computational savings and simplifications in the analysis of physical properties, motivating the development of cyclic- and helical-symmetry-adapted formulations in atomistic force-field methods James (2006); Dumitrica and James (2007); Dayal and James (2010); Aghaei and Dayal (2011); Aghaei et al. (2013), tight-binding approaches White et al. (1993); Allen (2007b, c); Popov (2004); Popov and Lambin (2006); Gunlycke et al. (2008); Zhang et al. (2009a, b); Kit et al. (2011); Zhang and Wei (2017), machine-learned force fields (MLFFs) Sharma et al. (2025), and Kohn–Sham DFT Saito et al. (1992); Ono and Hirose (2005a, b); Banerjee and Suryanarayana (2016); Ghosh et al. (2019); Banerjee (2021); Sharma and Suryanarayana (2021). In particular, the symmetry-adapted Kohn–Sham framework has enabled diverse applications, including bending moduli of 2D materials Kumar and Suryanarayana (2022, 2020), elastic moduli of nanotubes Bhardwaj et al. (2021); Bhardwaj and Suryanarayana (2022a), the flexoelectric effect in 2D materials Codony et al. (2021); Kumar et al. (2021), and impact of mechanical deformations on the electronic Bhardwaj and Suryanarayana (2023, 2022b) and spintronic Bhardwaj and Suryanarayana (2024) properties of nanotubes. However, these studies are restricted to static behavior and response.
@@ II Cyclic and helical symmetry
Consider a quasi-one-dimensional nanostructure comprising of $\hat{N}_{a}$ fundamental atoms and exhibiting cyclic and/or helical symmetry, as illustrated in Fig. 1. Let $\tilde{\Theta}$ denote the angle between an atom and its nearest cyclic image (Fig. 1(a)), and let $\varphi$ denote the circumferential-plane angle between an atom and its nearest helical image, which is separated by a distance $H$ along the longitudinal direction (Fig. 1(b)). The symmetry group of the nanostructure can be represented as a set of isometries obtained from the direct product (denoted by $\times$) of a cyclic symmetry group $\mathfrak{C}$ and a helical symmetry group $\mathfrak{H}$ McWeeny (2002):
@@
Indeed, the nanostructure is periodic along both the $\tilde{\theta}$ and $z$ directions in this helical coordinate system. Therefore, the helical coordinates of the image atom $\mathbf{\tilde{R}}_{I^{\prime}}$ are related to those of the corresponding fundamental atom $\mathbf{\tilde{R}}_{I}$ by a translation vector, i.e., $\mathbf{\tilde{R}}_{I^{\prime}}=\mathbf{\tilde{R}}_{I}+(0\,\,\zeta_{I^{\prime}}\tilde{\Theta}\,\,\mu_{I^{\prime}}H)^{T}$. Moreover, the 2D reciprocal space associated with the nanostructure in this coordinate system is discrete-continuous and periodic, with the first Brillouin zone (denoted by $\mathcal{U}$) defined by the set of $|\mathscr{G}|$ points $(\nu,\eta)$ such that, $\nu\in\{0,1,2,\cdots,\mathfrak{N}-1\}$ and $\eta\in[-\pi/H,\pi/H)$ Sharma and Suryanarayana (2021). These Brillouin zone points are commonly referred to as wavevectors. In what follows, the electron and phonon wavevectors are denoted by $\mathbf{k}=(0,\nu_{\mathbf{k}},\eta_{\mathbf{k}})$ and $\mathbf{q}=(0,\nu_{\mathbf{q}},\eta_{\mathbf{q}})$, respectively.
@@ III Symmetry-adapted DFT
The boundary condition satisfied by the symmetry-adapted Kohn–Sham orbitals in Eq. 12a corresponds to Bloch’s theorem in the helical coordinate system Sharma and Suryanarayana (2021). As noted in the previous section, the corresponding Brillouin zone consists of the points $(\nu,\eta)$ such that $\nu\in\{0,1,2,\cdots,\mathfrak{N}-1\}$ and $\eta\in[-\pi/H,\pi/H)$. The Hamiltonian depends on the electron density, which in turn depends on the orbitals, and therefore the governing equations must be solved self-consistently. A common approach is the self-consistent field (SCF) method, i.e., a fixed-point iteration in which the eigenproblem in Eq. 12a is solved to obtain the orbitals, from which, together with the occupations in Eq. 12c, the electron density is computed using Eq. 8. The electrostatic potential is then obtained by solving the Poisson problem in Eq. 12b; together with the exchange-correlation potential in Eq. 13a, it is used to update the eigenproblem, and the process is repeated until self-consistency is achieved. The resulting electronic ground-state quantities can then be used to evaluate the energy and its first-order derivatives, which includes the atomic forces and stress tensor Sharma and Suryanarayana (2021).
The cyclic- and helical-symmetry-adapted formalism presented above is for spin-unpolarized calculations within the pseudopotential approximation and at the LDA level of exchange and correlation. As in standard DFT, the pseudopotential approximation is adopted to reduce computational cost. The extension to all-electron calculations follows the standard DFT treatment Suryanarayana et al. (2011, 2010): the nonlocal pseudopotential energy term in Eq. 7 and the corresponding operator in the Hamiltonian in Eq. 12a are set to zero, and the ionic pseudopotential is replaced by the corresponding Coulomb potential. The framework can also be generalized to collinear spin-polarized calculations by solving separate eigenproblems for the two spin channels, which are coupled through the spin-resolved and total electron densities, following standard DFT. It can further be extended to semilocal exchange-correlation functionals, as demonstrated in earlier cyclic- and helical-symmetry-adapted work Sharma and Suryanarayana (2021).
@@ IV.1 Perturbative expansion
The atomic displacement in the Cartesian coordinate system can be obtained using the transformation:
@@ IV.2 Sternheimer equation
where $\mathcal{I}$ denotes the identity operator, and the coefficients $\xi_{m,\mathbf{k}+\mathbf{q}}$ are chosen to remove the singularity and improve the conditioning of the equation, thereby ensuring robust and faster convergence to the solution.
The cyclic- and helical-symmetry-adapted DFPT formalism presented above can be extended to all-electron calculations by replacing the ionic pseudopotential with the Coulomb potential and setting both the nonlocal pseudopotential operator in the Hamiltonian $\mathcal{H}_{\mathbf{k}+\mathbf{q}}$ and its perturbation ${\mathcal{V}_{nl}}^{(1)}_{\mathbf{k},\mathbf{q},J_{\beta}}$ in Eq. 21 to zero. It can also be generalized to spin-polarized systems by solving the Sternheimer equations separately for the orbital perturbations in each spin channel, with the two channels coupled through the perturbations in the spin-resolved and total electron densities, following standard DFPT Dal Corso (2001). The extension to semilocal exchange-correlation functionals requires a corresponding modification of ${\mathcal{V}_{xc}}^{(1)}_{\mathbf{q},J_{\beta}}$ in Eq. 21, as in standard DFPT Dal Corso and de Gironcoli (2000).
@@ IV.3 Dynamical matrix
The cyclic- and helical-symmetry-adapted expression for the dynamical matrix presented above is directly applicable to semilocal exchange-correlation functionals. Within the pseudopotential approximation, its extension to spin-polarized systems requires summation over the two spin channels for the nonlocal pseudopotential contributions, namely, the last three terms in Eq. 28. As noted earlier, the all-electron formulation is obtained by replacing the ionic pseudopotential with the Coulomb potential and setting the nonlocal pseudopotential contributions to zero.
@@ V Implementation
We have implemented the cyclic- and helical-symmetry-adapted Kohn-Sham DFT phonon framework within the Cyclix-DFT Sharma and Suryanarayana (2021) feature of the real-space code M-SPARC Xu et al. (2020); Zhang et al. (2023), which is a Matlab-based variant of the large-scale electronic-structure code SPARC Xu et al. (2021); Zhang et al. (2024). In particular, all quantities in the Kohn-Sham formalism are represented on a uniform grid in the aforementioned helical coordinate system. The differential operators are approximated using high order centered finite-differences and spatial integrations are approximated using the trapezoidal rule. The Brillouin zone is discretized at the points $(\nu,\eta_{MP}+\frac{\varphi}{H}\nu)$, where $\eta_{MP}$ is generated using the Monkhorst-Pack Monkhorst and Pack (1976) scheme. The Perdew-Zunger Perdew and Zunger (1981) variant of the LDA Kohn and Sham (1965) exchange-correlation, optimized norm-conserving Vanderbilt pseudopotentials Hamann (2013), and Fermi-Dirac electronic smearing are employed.
@@ VI Results and discussion
where $a=\sqrt{3}\,a_{0}$ is the length of graphene’s lattice vector and $r$ is the nanotube radius. This scaling law is in very good qualitative agreement with previous studies Sharma et al. (2025); Gunlycke et al. (2008). The prefactor of $45~\mathrm{cm}^{-1}$ obtained here is is very good agreement with the value of $46~\mathrm{cm}^{-1}$ reported recently using a cyclic- and helical-symmetry-adapted machine-learned force field Sharma et al. (2025), but differs noticeably from the value of $54~\mathrm{cm}^{-1}$ obtained earlier using an interatomic potential Gunlycke et al. (2008).
@@
This scaling law is also in very good qualitative agreement with the literature Sharma et al. (2025); Gunlycke et al. (2008). However, the constant of $468$ cm-1 obtained here is noticeably different from the values of $480$ cm-1 and $486$ cm-1 obtained previously using machine-learned force fields (MLFFs) Sharma et al. (2025) and interatomic potential Gunlycke et al. (2008), respectively.
@@ VII Concluding remarks
In this work, we developed a first-principles framework for the calculation of phonons in nanostructures with cyclic and/or helical symmetry. In particular, we derived cyclic- and helical-symmetry-adapted representations for the dynamical matrix at arbitrary phonon wavevectors within a variationally formulated, symmetry-adapted DFPT framework. In addition, we derived the acoustic sum rules for cylindrical geometries, which include a rigid-body rotational mode in addition to the three translational modes. The resulting formalism was implemented within a high-order finite-difference discretization. Using carbon nanotubes as representative systems, we validated the accuracy of the framework through excellent agreement with periodic plane-wave results. We further applied the framework to compute the Young’s and shear moduli of carbon nanotubes, as well as the scaling laws governing the dependence of ring and radial breathing mode phonon frequencies on nanotube diameter. The elastic moduli were found to be in agreement with previous DFT and experimental results, while the phonon scaling laws showed qualitative agreement with previous atomistic simulations. Overall, the symmetry-adapted formalism enables a substantial reduction in computational cost for systems with cyclic and/or helical symmetry. For example, in the case of a (16,0) carbon nanotube, the symmetry-adapted formulation requires only 2 atoms in the fundamental domain, compared to 64 atoms under standard periodic boundary conditions. These advantages become even more pronounced for chiral nanotubes, with increasing diameter, and under applied twist, where the reduction in the number of atoms and the associated computational effort is significantly greater.
@@ ACKNOWLEDGMENTS
This work draws from the the thesis of A.S. at Georgia Institute of Technology. A.S. gratefully acknowledges the ANRF early career research grant (Grant No. ANRF/ECRG/2024/002362) from Department of Science and Technology India. A.S. and P.S. gratefully acknowledge support from the U.S. National Science Foundation through CAREER Grant No. 155321, during the time A.S. was at Georgia Institute of Technology. This research was also supported by the supercomputing infrastructure provided by Partnership for an Advanced Computing Environment (PACE) through its Phoenix cluster at Georgia Institute of Technology, Atlanta, Georgia.
Data availability statements
The data that support the findings of this article are openly available Sharma and Suryanarayana (2026).